Barcode Biz has invested in new machinery at a cost of $1,450,000. This investment is expected to produce cash flows of $640,000, $715,250, $823,330, and $907,125 over the next four years. What is the payback period for this project (rounded to two decimal places)? Select one: A. 2.12 years B. 1.88 years C. 4.00 years D. 3.00 years.
Answer #1

Cumulative cash flows till year 0=-1450000
Cumulative cash flows till year 1=-1450000+640000=-810000
Cumulative cash flows till year 2=-1450000+640000+715250=-94750
Cumulative cash flows till year 3=-1450000+640000+715250+823330=728580

Hence, payback period=2+94750/823330=2.115081438 years